# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ 4 Optimal approximations of covariance operators
In this section, we formulate a minimisation problem that aims at finding low-rank approximations of $\mathcal{C}_{\textup{pos}}$ that are optimal simultaneously with respect to all members of a class of spectral loss functions. This class includes the forward and reverse Kullback–Leibler divergence and the Rényi divergences as special cases.
4.1 Spectral loss functions and approximation classes
To measure the quality of a given approximation of the exact posterior covariance $\mathcal{C}_{\textup{pos}}$, we define a class of loss functions on $\mathcal{E}^{2}$ in the following way. Define
@@ 4.1 Spectral loss functions and approximation classes
$\mathscr{L}$×$\coloneqq\left\{\mathcal{E}\times\mathcal{E}\ni(\mathcal{C}_{2},\mathcal{C}_{1})\mapsto\mathcal{L}_{f}(\mathcal{C}_{2}\|\mathcal{C}_{1})\coloneqq\sum_{i}f(\Lambda_{i}(R(\mathcal{C}_{2}\|\mathcal{C}_{1})))\right\}.$ (15b)
As we show in Lemma 4.1 below, the conditions $f\in C^{1}$ and $xf^{\prime}(x)>0$ for $x\not=0$ ensure that 00 is the unique minimiser of every $f\in\mathscr{F}$. The Lipshitz continuity of $f^{\prime}$ at 00 implies that every $\mathcal{L}\in\mathscr{L}$ is takes only finite values and, as will be shown later in Lemma 4.9, that $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\cdot)$ is differentiable on a suitable subspace of $\mathcal{E}$. The blowup at infinity condition is used to prove coercivity of $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\cdot)$ on a suitable subspace of $\mathcal{E}$, as we show in Lemma 4.16.
@@
Let $f\in\mathscr{F}$. Then
@@
    Let $\alpha:(-1,\infty)\rightarrow(-1,\infty)$ be defined by $\alpha(x)=\frac{-x}{1+x}$. If $f\in\mathscr{F}$ satisfies $\lim_{x\rightarrow-1}f(x)=\infty$, then $f\circ\alpha\in\mathscr{F}$.
The class of loss functions considered in the finite-dimensional of [30, Definition 2.1] differs from the class (15) in two aspects. First, the domain of the former functions $f$ is $(0,\infty)$ and $f(0)$ need not be 0, while the domain of the latter is $(-1,\infty)$ and we require $f(0)=0$. That the natural class to consider involves the horizontal shift of $-1$ and the vertical shift becomes apparent as the fundamental object governing the losses is given by (9), which is a compact operator and therefore has an eigenvalue sequence accumulating at 0. Second, there is an additional Lipschitz condition in (15a), which implies that $\mathcal{L}$ is finite on $\mathcal{E}^{2}$ for every $\mathcal{L}\in\mathscr{L}$. Note that the condition that $f^{\prime}$ is Lipschitz continuous at 00 is not implied by the other conditions in (15a), as the function $f$ with $f^{\prime}(x)=sgn(x)\lvert x\rvert^{\alpha}$, $\alpha\in(0,1)$, and $f(0)=0$ shows. Here, $sgn(x)$ denotes the function assigning $1$ to $x\geq 0$ and $-1$ otherwise. This function satisfies all conditions of (15a) except the Lipschitz condition of $f^{\prime}$ at 0.
While restricted compared to [30], the loss class (15) is still rich enough to include the forward and reverse KL divergences and Réyni divergence between equivalent Gaussian measures with the same mean. This partially extends [30, Lemma 2.2], in which this is shown for the forward KL and Hellinger distance in the finite-dimensional setting.
@@
Let $m\in\mathcal{H}$. Let $\mu_{i}=N(m_{i},\mathcal{C}_{i})$ be nondegenerate and $\mathcal{C}_{i}\in\mathcal{E}$ for $i=1,2$.
@@
    For the reverse distances, we have $f_{\textup{KL}}\circ\alpha,f_{\textup{Ren},\rho}\circ\alpha\in\mathscr{F}$ with $\alpha(x)=\frac{-x}{1+x}$ on $(-1,\infty)$, and
    $D_{\textup{KL}}(\mu_{1}\|\mu_{2})=\mathcal{L}_{f_{\textup{KL}}\circ\alpha}(\mathcal{C}_{2}\|\mathcal{C}_{1}),\quad D_{\textup{Ren},\rho}(\mu_{1}\|\mu_{2})=\mathcal{L}_{f_{\textup{Ren},\rho}\circ\alpha}(\mathcal{C}_{2}\|\mathcal{C}_{1}).$
Given the approximation classes (4) and (5) and given the covariance loss functions in (15b), we can define the following low-rank approximation problem, for every $r\leq n$.
@@
We note that even if an optimal covariance and precision can be found for given $\mathcal{L}\in\mathscr{L}$, it is not a priori clear that they are in fact independent of $\mathcal{L}$.
@@
We also emphasise the following. Since the inverse of a self-adjoint positive matrix is again self-adjoint and positive, inverses of covariance operators are covariance operators in the case $\dim{\mathcal{H}}<\infty$. If $\dim{H}=\infty$, then this no longer holds because the trace-class property is not preserved under inversion in infinite dimensions. In fact, if $\dim{H}=\infty$ and $T$ is trace class, then $T^{-1}$ is an unbounded operator and its eigenvalue sequence is not summable since it is not bounded. In order to define a loss on precisions, as is done in the finite-dimensional case of [30, Corollary 3.1], one approach is to extend the domain of $\mathcal{L}\in\mathscr{L}$ to $\mathcal{E}^{2}\cup(\mathcal{E}^{-1})^{2}$ via $\mathcal{L}_{ext}(\mathcal{C}_{1}^{-1}\|\mathcal{C}_{2}^{-1})\coloneqq\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})$ and $\mathcal{L}_{ext}(\mathcal{C}_{1}\|\mathcal{C}_{2})\coloneqq\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})$ for $\mathcal{C}_{1},\mathcal{C}_{2}\in\mathcal{E}$. If $(\mathcal{C}_{1},\mathcal{C}_{2})\in\mathcal{E}^{2}\cap(\mathcal{E}^{-1})^{2}$, then $\mathcal{L}_{ext}(\mathcal{C}_{1}^{-1}\|\mathcal{C}_{2}^{-1})=\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})=\mathcal{L}_{ext}(\mathcal{C}_{1}\|\mathcal{C}_{2})$, showing that $\mathcal{L}_{ext}$ is well-defined. By the definitions (15), (9) and Lemma 3.3, $\mathcal{L}$ depends on $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ only via the set of eigenvalues of the extensions of the densely defined and bounded operator $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}\mathcal{C}_{1}^{-1/2}-I$. Lemma 3.3(i)-(ii) shows that eigenvalues of the latter operator remain unchanged when replacing $\mathcal{C}_{1}$ by $\mathcal{C}_{2}^{-1}$ and $\mathcal{C}_{2}$ by $\mathcal{C}_{1}^{-1}$. Thus, $\mathcal{L}_{ext}(\mathcal{C}_{2}^{-1}\|\mathcal{C}_{1}^{-1})=\mathcal{L}_{ext}(\mathcal{C}_{1}\|\mathcal{C}_{2})=\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})$. This equation firstly implies that $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{P}^{-1})=\mathcal{L}_{ext}(\mathcal{P}\|\mathcal{C}_{\textup{pos}}^{-1})$ for $\mathcal{P}\in\mathscr{P}_{r}$, and we can reformulate 4.4 accordingly in terms of a loss $\mathcal{L}_{ext}$ on precisions. Secondly, it shows that there is no need to explicitly define a loss on precisions, as we can just use $\mathcal{L}$ on the corresponding covariances in reversed order instead.
A natural question is what the precisions corresponding to approximate covariances in $\mathscr{C}_{r}$ are, and what the covariances corresponding approximate precisions in $\mathscr{P}_{r}$ are. The next result shows that $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$ are in fact each others inverses and that Problems 4.3 and 4.4 are in fact equivalent. Item (i) can be seen as a generalisation of the finite-dimensional result [30, Lemma A.2] to infinite-dimensional Hilbert spaces.
@@
Let $r\in\mathbb{N}$.
@@
    For every $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ such that $\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathscr{C}_{r}$, there exists $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the same rank as $K$, such that $(\mathcal{C}_{\textup{pr}}-KK^{*})^{-1}=\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\in\mathscr{P}_{r}$. The reverse correspondence also holds: for every $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ such that $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\in\mathscr{P}_{r}$, there exists $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the same rank as $U$, such that $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}=\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathscr{C}_{r}$. In particular, $\mathscr{C}_{r}^{-1}\coloneqq\{\mathcal{C}^{-1}:\ \mathcal{C}\in\mathscr{C}_{r}\}=\mathscr{P}_{r}$ and $\mathscr{P}_{r}^{-1}\coloneqq\{\mathcal{P}^{-1}:\ \mathcal{P}\in\mathscr{P}_{r}\}=\mathscr{C}_{r}$

### v3

@@ 4 Optimal approximations of covariance operators
In this section, we formulate a minimisation problem that aims at finding low-rank approximations of $\mathcal{C}_{\textup{pos}}$ that are optimal simultaneously with respect to all members of a class of spectral loss functions. This class includes the Rényi divergences and forward and reverse KL divergences as special cases. The loss class and the low-rank covariance approximation problems are introduced in Section 4.1, the equivalence to the exact posterior of the approximations considered in Section 2 is studied in Section 4.2, the approximation problems are formulated as minimisation problems involving a differentiable function in Section 4.3, and the approximation problems are solved in Section 4.4. The proofs of all the results in this section are given in Section B.2.
4.1 Spectral loss functions and problem formulation
To measure the quality of a given approximation of the exact posterior covariance $\mathcal{C}_{\textup{pos}}$, we define a class of loss functions on $\mathcal{E}^{2}$ in the following way. Recall the definition of the eigenvalue map $\Lambda$ defined on Hilbert–Schmidt operators, from Section 1.5. Also recall the definition of the Hilbert–Schmidt operator-valued map $R(\cdot\|\cdot)$ from (9). Define
@@ 4.1 Spectral loss functions and problem formulation
$\mathscr{L}$×$\coloneqq\left\{\mathcal{E}\times\mathcal{E}\ni(\mathcal{C}_{2},\mathcal{C}_{1})\mapsto\mathcal{L}_{f}(\mathcal{C}_{2}\|\mathcal{C}_{1})\coloneqq\sum_{i}f(\Lambda_{i}(R(\mathcal{C}_{2}\|\mathcal{C}_{1}))):\ f\in\mathscr{F}\right\}.$ (16b)
As we show in Lemma 4.1 below, the conditions $f\in C^{1}((-1,\infty))$ and $xf^{\prime}(x)>0$ for $x\not=0$ ensure that 00 is the unique minimiser of every $f\in\mathscr{F}$. The Lipschitz continuity of $f^{\prime}$ at 00 implies that $(f(x_{i}))_{i}$ is summable for $(x_{i})_{i}\in\ell^{2}((-1,\infty))$, so that every $\mathcal{L}\in\mathscr{L}$ takes only finite values. Furthermore, this Lipschitz continuity implies that $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\cdot)$ is differentiable on a suitable subspace of $\mathcal{E}$, as will be shown later in Lemma 4.11. The blowup at infinity condition is used to prove coercivity of $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\cdot)$ on suitable subspaces of $\mathcal{E}$, as we show in Lemma 4.19.
@@
Let $\mathscr{F}$ be given in (16a) and $f\in\mathscr{F}$. Then
@@
    Let $\eta:(-1,\infty)\rightarrow(-1,\infty)$ be defined by $\eta(x)=\frac{-x}{1+x}$. If $f\in\mathscr{F}$ satisfies $\lim_{x\rightarrow-1}f(x)=\infty$, then $f\circ\eta\in\mathscr{F}$.
The class of loss functions considered in the finite-dimensional setting of [31, Definition 2.1] differs from the class (16) in two aspects. For every function $f$ in the former, the domain is $(0,\infty)$ and $f$ need not have minimum equal to 0, while for every function in the latter, the domain is $(-1,\infty)$ and we require $f(0)=0$. That the natural class to consider involves the horizontal shift of $-1$ and the vertical shift becomes apparent as the fundamental object governing the losses is given by the operator $R(\mathcal{C}_{2}\|\mathcal{C}_{1})$ defined in (9), which is a compact operator and therefore has an eigenvalue sequence accumulating at 0. Second, there is an additional Lipschitz condition in (16a), which implies that $\mathcal{L}$ is finite on $\mathcal{E}^{2}$ for every $\mathcal{L}\in\mathscr{L}$. Note that the condition that $f^{\prime}$ is Lipschitz continuous at 00 is not implied by the other conditions in (16a), as the function $f$ with $f^{\prime}(x)=\operatorname{sgn}(x)\lvert x\rvert^{\alpha}$, $\alpha\in(0,1)$, and $f(0)=0$ shows. Here, $\operatorname{sgn}(x)$ denotes the function assigning $1$ to $x\geq 0$ and $-1$ otherwise. This function satisfies all conditions of (16a) except the Lipschitz condition of $f^{\prime}$ at 0.
While restricted compared to the class in [31, Definition 2.1], the class (16) is still rich enough to include the forward and reverse KL divergences and Rényi divergences between equivalent Gaussian measures with the same mean, as shown in the following result. This result partially extends [31, Lemma 2.2], in which the analogous statement is shown for the forward KL divergence in the finite-dimensional setting.
@@
Let $m\in\mathcal{H}$. Let $\mu_{i}=\mathcal{N}(m,\mathcal{C}_{i})$ be nondegenerate and $\mathcal{C}_{i}\in\mathcal{E}$ for $i=1,2$.
@@
    For the reverse divergences, we have $f_{\textup{KL}}\circ\eta,f_{\textup{Ren},\rho}\circ\eta\in\mathscr{F}$ with $\eta(x)\coloneqq\frac{-x}{1+x}$ on $(-1,\infty)$, and
    $D_{\textup{KL}}(\mu_{1}\|\mu_{2})=\mathcal{L}_{f_{\textup{KL}}\circ\eta}(\mathcal{C}_{2}\|\mathcal{C}_{1}),\quad D_{\textup{Ren},\rho}(\mu_{1}\|\mu_{2})=\mathcal{L}_{f_{\textup{Ren},\rho}\circ\eta}(\mathcal{C}_{2}\|\mathcal{C}_{1}).$
Given the approximation classes (4) and (5) and given the covariance loss functions in (16b), we can define the following low-rank approximation problem, for every $r\leq n$. We do not consider the case $r>n$, because in this case the problems have the trivial solutions $\mathcal{C}_{\textup{pos}}$ and $\mathcal{C}_{\textup{pos}}^{-1}$ respectively.
@@
We note that even if an optimal covariance and precision can be found for some given $\mathcal{L}\in\mathscr{L}$, it is not a priori clear that they are in fact independent of $\mathcal{L}$.
@@
We also emphasise the following. Since the inverse of a self-adjoint positive matrix is again self-adjoint and positive, inverses of covariance operators are covariance operators in the case $\dim\mathcal{H}<\infty$, but not if $\dim\mathcal{H}=\infty$. This is because the trace-class property is not preserved under inversion in infinite dimensions. In fact, if $\dim{\mathcal{H}}=\infty$ and $T$ is trace class, then $T^{-1}$ is an unbounded operator and its eigenvalue sequence is not summable since this sequence is not bounded. In order to define a loss on precisions, as is done in the finite-dimensional case of [31, Corollary 3.1], one approach is to extend the domain of $\mathcal{L}\in\mathscr{L}$ to $\mathcal{E}^{2}\cup(\mathcal{E}^{-1})^{2}$ via $\mathcal{L}_{ext}(\mathcal{C}_{1}^{-1}\|\mathcal{C}_{2}^{-1})\coloneqq\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})$ and $\mathcal{L}_{ext}(\mathcal{C}_{1}\|\mathcal{C}_{2})\coloneqq\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})$ for $\mathcal{C}_{1},\mathcal{C}_{2}\in\mathcal{E}$. If $(\mathcal{C}_{1},\mathcal{C}_{2})\in\mathcal{E}^{2}\cap(\mathcal{E}^{-1})^{2}$, then $\mathcal{L}_{ext}(\mathcal{C}_{1}^{-1}\|\mathcal{C}_{2}^{-1})=\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})=\mathcal{L}_{ext}(\mathcal{C}_{1}\|\mathcal{C}_{2})$, showing that $\mathcal{L}_{ext}$ is well-defined. By the definitions (16), (9) and Lemma 3.4, $\mathcal{L}$ depends on $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ only via the set of eigenvalues of the bounded extensions of the densely defined operator $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}\mathcal{C}_{1}^{-1/2}-I$. Lemma 3.4(i)-(ii) show that eigenvalues of the latter operator remain unchanged when replacing $\mathcal{C}_{1}$ by $\mathcal{C}_{2}^{-1}$ and $\mathcal{C}_{2}$ by $\mathcal{C}_{1}^{-1}$. Thus, $\mathcal{L}_{ext}(\mathcal{C}_{2}^{-1}\|\mathcal{C}_{1}^{-1})=\mathcal{L}_{ext}(\mathcal{C}_{1}\|\mathcal{C}_{2})=\mathcal{L}(\mathcal{C}_{1}\|\mathcal{C}_{2})$. This equation firstly implies that $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{P}^{-1})=\mathcal{L}_{ext}(\mathcal{P}\|\mathcal{C}_{\textup{pos}}^{-1})$ for $\mathcal{P}\in\mathscr{P}_{r}$, and we can reformulate Problem 4.4 accordingly in terms of a loss $\mathcal{L}_{ext}$ on precisions. Secondly, it shows that there is no need to explicitly define a loss on precisions, as we can just use $\mathcal{L}$ on the corresponding covariances in reverse order instead.
4.2 Equivalence to target measures of low-rank Gaussian approximations
As discussed in Section 3, not all approximations $\mathcal{N}(m_{\textup{pos}},\mathcal{C}_{\textup{pr}}-KK^{*})$ are probability measures equivalent to $\mu_{\textup{pos}}$. This equivalence holds only if $\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathcal{E}$, with $\mathcal{E}$ defined in (6). The first aim of this section is to characterise the sets $\mathscr{C}_{r,\mathcal{E}}\coloneqq\{\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathcal{E}:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$ and $\mathscr{P}_{r,\mathcal{E}}\coloneqq\{(\mathcal{C}_{\textup{pr}}-KK^{*})^{-1}:\ \mathcal{C}_{\textup{pr}}-KK^{*}\in\mathcal{E},\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$, which is done in Lemma 4.5(iii) and Proposition 4.6(i) respectively. We write $\mathscr{C}_{r,\mathcal{E}}^{-1}\coloneqq\{\mathcal{C}^{-1}:\ \mathcal{C}\in\mathscr{C}_{r,\mathcal{E}}\}=\mathscr{P}_{r,\mathcal{E}}$. The results are formulated for arbitrary Gaussian measures, because they are not intrinsic to the Bayesian formulation. We also show that $\mathscr{C}_{r}\subset\mathscr{C}_{r,\mathcal{E}}$ and $\mathscr{P}_{r}\subset\mathscr{P}_{r,\mathcal{E}}$, with $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$ from (4) and (5) respectively, and that these inclusions are strict. In this section, we also determine the relationship between $\mathscr{C}_{r}^{-1}\coloneqq\{\mathcal{C}^{-1}:\ \mathcal{C}\in\mathscr{C}_{r}\}$ and $\mathscr{P}_{r}$, and between Problem 4.3 and Problem 4.4.
We shall characterise the elements in $\mathcal{E}$ of the form $\mathcal{C}_{\textup{pr}}-KK^{*}$ with $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ for some $r\in\mathbb{N}$ using Lemma 4.5. Because this result is not intrinsic to the Bayesian interpretation, we formulate it for the more general set $\mathcal{E}(m_{1},\mathcal{C}_{1})$ defined in (7), which contains all covariances $\mathcal{C}$ such that $\mathcal{N}(m_{1},\mathcal{C})\sim\mathcal{N}(m_{1},\mathcal{C}_{1})$, for arbitrary Gaussian target measures $\mathcal{N}(m_{1},\mathcal{C}_{1})$ with $m_{1}\in\mathcal{H}$ and injective $\mathcal{C}_{1}\in L_{1}(\mathcal{H})_{\mathbb{R}}$. Lemma 4.5 shows that the operator $I-(\mathcal{C}_{1}^{-1/2}K)(\mathcal{C}_{1}^{-1/2}K)^{*}$ is important for determining whether $\mathcal{C}=\mathcal{C}_{1}-KK^{*}$ belongs to $\mathcal{E}(m_{1},\mathcal{C}_{1})$. Item (ii) shows that the assumption that $I-(\mathcal{C}_{1}^{-1/2}K)(\mathcal{C}_{1}^{-1/2}K)^{*}$ is well-defined and nonnegative, is equivalent to $\mathcal{C}\geq 0$, which is necessary for $\mathcal{C}\in\mathcal{E}(m_{1},\mathcal{C}_{1})$. Item (iii) shows that this assumption with the additional assumption of invertibility of $I-(\mathcal{C}_{1}^{-1/2}K)(\mathcal{C}_{1}^{-1/2}K)^{*}$ is both necessary and sufficient for $\mathcal{C}\in\mathcal{E}(m_{1},\mathcal{C}_{1})$. By item (i), $I-(\mathcal{C}_{1}^{-1/2}K)(\mathcal{C}_{1}^{-1/2}K)^{*}$ is well-defined under the range condition $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$. Item (iv) relates the properties of the eigenvalues of $I-(\mathcal{C}_{1}^{-1/2}K)(\mathcal{C}_{1}^{-1/2}K)^{*}$ to the properties $\mathcal{C}\geq 0$ or $\mathcal{C}\in\mathcal{E}(m_{1},\mathcal{C}_{1})$ of $\mathcal{C}$. If $\mathcal{C}>0$, then item (iv) together with Lemma A.1 also shows that the range condition $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}}$ implies the diagonalisability of $I-(\mathcal{C}_{1}^{-1/2}K)(\mathcal{C}_{1}^{-1/2}K)^{*}$ in the Cameron–Martin space of $\mathcal{N}(m_{1},\mathcal{C}_{1})$.
@@ 4.2 Equivalence to target measures of low-rank Gaussian approximations
Let $\mathcal{C}_{1}\in L_{1}(\mathcal{H})_{\mathbb{R}}$ be injective and $m_{1}\in\mathcal{H}$. Let $\mathcal{C}\coloneqq\mathcal{C}_{1}-KK^{*}$ for some $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ and $r\in\mathbb{N}$. The following holds:
@@
    If $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$, then $\mathfrak{X}\coloneqq I-(\mathcal{C}_{1}^{-1/2}K)(\mathcal{C}_{1}^{-1/2}K)^{*}$ is well-defined and $\mathcal{C}=\mathcal{C}_{1}^{1/2}\mathfrak{X}\mathcal{C}_{1}^{1/2}$,